b) The quarterly demand for an item of raw materials is estimated at 2,000 units at a purchase price of GH¢180 per unit. It is estimated that the cost per order will be GH¢270 and the cost of holding a unit of material in inventory will be GH¢24. Required:
i. Compute the optimal order quantity, and total minimum costs. [5 marks] ii. Suppose a supplier offers 5% quantity discount for purchase of 8,000 units, should the offer be accepted [4 marks]
c) State ten (10) activities likely to be undertaken by cost and management accounting department of a manufacturing company [10 marks]
d
Answer #1
EOQ
Details No /Amt GHC
Annual usgae units =2000 qtrly *4=                                                      8,000
Cost per Order $                                                     270
Carrying cost per unit = $                                                 24.00
EOQ = Sq Root of [(2*Order cost*Annual usage)/(Annual Holding cost per unit]
=Sq Root (2*270*8000/24)
EOQ =424.264 units
So the optimal order qty is 424.264 units.
Total no of orders in a year=8000/424.264=18.86.
Total Ordering cost=18.86*270=GHC 5091.17
Average inventory=424.264/2=212.13
Inventory holding cost =212.13*24=5091.17
So total minimum cost =5091.17*2=10,182.34 Ans i
Ans ii)
Assume the order Qty =8000 units
Average inventory =8000/2=4000 units
Inventory Holding cost =4000*24=96,000
Ordering cost =270
Total ordering and holding cost =96,270
Incremental cost over Minimal ordering=96270-10182.34=86,087.7
Savings in Purchase Price @5%=8000*270*5%=108,000
Incremental savings over optimal order=108000-86087.7=21,912.34
Therefore , the quantity discount should be accepted. Ans ii
Ans c
Ten Activities that may linkely be undertaken bu cost & Management accounting Department are ;
1. Preparation of Budget for the company
2. Doing Budget vs Actual Analysis
3. Setting Standard cost of products and services
4. Doing Marginal costing and contribution Margin Analysis
5. Doing Process order/Job Order costing
6. Performing Activity Based costing
7. Doing evaluation of Capital Investment Options.
8. Doing analysis for make or buy decisions.
9. Doing Ratio Analysis and benchmarking of performance.
10. Doing Product wise profitability analysis.